• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Ole Hansen

    Log entry time 15:57:41 on March28,2005

    Entry number 142641

    keyword=Shift Summary

    8:00 Start shift with run 3017 in progress.
    8:30 Beam off for beam studies
    9:05 Q1 trip. Cannot be reset from upstairs.
    10:00 Q1 reset and ramping up. Hall in Controlled Access.
    11:57 Another Q1 trip. Again caused by a drop in He cryo level. Manage to reset from upstairs.
    13:15 Hall closed and in Beam Permit. Beam studies still ongoing.
    14:30 Beam studies completed. No beam because of IOC problems.
    16:00 Still no beam. Apparently, there is no IOC problem, but rather a site-wide problem with a synchronization signal.


    Runs
    ----
    3017 - 161k - good run, continued from owl shift